The perimeter of a window is 32 feet. The width is 6 times bigger than the length. What is the width?
Answer by Alan3354(69443)

The perimeter of a window is 32 feet. The width is 6 times bigger than the length. What is the width?
-------
P = 2W + 2L = 32
W + L = 16
-------
W = 7L
----
7L + L = 16
L = 2 feet
W = 14 feet